News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 26 August 2022 9.00am EDT

During the last week the Chandra observing schedule was interrupted at 2:36pm EDT on Aug 19 by what was likely a single event upset (SEU) in the telemetry processor of the Command and Telemetry Unit, causing a reset of its timing signals and resets of several spacecraft components. An elevated temperature of the aspect camera CCD due to the resets led to its losing track of stars resulting in the spacecraft entering normal-sun mode. All safing actions were nominal.

Recovery activities were successfully executed Aug 19-20 following standard operating procedures for recovering from a CTU reset and recovering from Normal Sun Mode. The OBC-A and -B memory were dumped as a follow-up to the flight software patch. The dumps will be used to update the baseline memory images maintained on the ground.

In addition to the real-time procedures associated with the CTU reset recovery, real-time procedures were executed on Aug 19 and Aug 22 to initiate long ACIS CTI data collection while waiting to return to observing science.

A new mission schedule was uplinked on Aug 22, which resumed science observations at 10:32am EDT on Aug 23. There were 257.0 ks of science loss due to the anomaly. Planned observations were impacted by the interruption and will be rescheduled.

Another real-time procedure was executed on Aug 25 to check out HRC-I operation on the A-side electronics. As part of the continued evaluation of the HRC, the HRC 15V power supply was successfully powered on for approximately 10 ks and an HRC-I science observation of TYC 6991-475-1 was completed.

The schedule of targets for the next week is shown below.

        GDH8                   ACIS-I       Aug 28
        GDH8                   ACIS-I
        GDH8                   ACIS-I
        PSZ2G026.73-38.94      ACIS-I
        3c297                  ACIS-S       Aug 29
        WINGSJ211347.41+022    ACIS-S
        3c297                  ACIS-S
        GDH8                   ACIS-I
        GDH8                   ACIS-I
        GDHIdp55               ACIS-I
        Radiation Belts                     Aug 30
        H1413+117              ACIS-S
        CMZMolecularCloud      ACIS-I
        WINGSJ211347.41+022    ACIS-S
        GDH25 (4 obs)          ACIS-I
        H1413+117              ACIS-S       Aug 31
        CFHQSJ210054-171522    ACIS-S
        GDH25                  ACIS-I
        GDHIdp55 (5 obs)       ACIS-I
        CFHQSJ210054-171522    ACIS-S
        NGC5759                ACIS-S       Sep  1
        SDSSJ093138.51+5633    ACIS-S
        Radiation Belts
        Abell2744              ACIS-I
        NGC6357field2          ACIS-I       Sep  2
        Abell2744              ACIS-I
        2RXSJ141916.4+13001    ACIS-S
        PSZ2G073.97-27.82      ACIS-I
        Abell2142              ACIS-I
        Abell2744              ACIS-I       Sep  3
        CMZMolecularCloud      ACIS-I
        Abell2744              ACIS-I
        Radiation Belts                     Sep  4
        Abell2744              ACIS-I
        Abell2142              ACIS-I


All spacecraft subsystems except HRC continued to support nominal operations.

Return to Status Reports